OpenStudy (anonymous):

a man walks 6 miles at 4 miles per hour. At what speed would he need to travel during the next 2 1/2 hours to have an average speed of 6 miles per hour during the complete trip?

OpenStudy (anonymous):

Avg speed = distance / time, right?

OpenStudy (anonymous):

So first, for how long will he have walked in total?

OpenStudy (anonymous):

He walks 6 miles at 4 mi/hr, so 6mi / 4mi/hr = 3/2 hrs. Then he walks for 2 1/2, or 5/2 hrs. So in total he walks for 8/2 or 4 hrs.

OpenStudy (anonymous):

He walks for 4 hours and wants to average 6 mi/hr, so he needs to have a total of 6*4 = 24 miles covered. He has walked 6 already, so he needs to walk 18 in the next 2 1/2 hours. 18 / (5/2) = 36/5 mi/hr

OpenStudy (anonymous):

Kablam, sucka.

OpenStudy (anonymous):

thank you